Logiweb bibliography of check

Up Help

[0] check

[1] base

The pyk compiler, version 0.grue.20060417+ by Klaus Grue GRD-2006-05-04.UTC:15:42:31.918490 = MJD-53859.TAI:15:43:04.918490 = LGT-4653474184918490e-6